# Structured Learning of Compositional Sequential Interventions

### v1

@@ 2.1 Representation Power and Compositional Warping
Eq. (1) is reminiscent of tensor factorization and its uses in causal modeling (Athey et al., 2021; Agarwal et al., 2023b), where the primary motivation comes from the imputation of missing potential outcomes taking place on a pre-determined period in the past. From a predictive perspective, it can be motivated from known results in functional analysis (such as Proposition 1 of (Kaddour et al., 2021)) that allows us to represent a function $f(x)$ by first partitioning its input into $(x_{a},x_{b})$ and controlling the approximation error given by an inner product of vector-valued functions $f(x_{a},x_{b})\approx f_{a}^{\mathsf{T}}(x_{a})f_{b}(x_{b})$, in practice choosing the dimensionality of the vectors by a data-driven approach. Although those results apply typically to smooth functions with continuous inputs (such as the Taylor series approximation, which relies on vectors of monomials with coefficients constructed from derivatives), discrete inputs such as e.g. $D_{n}^{1:T}$ at a fixed dimensionality $T$ can be separated from the other inputs as $f(n,d^{1:T},x^{1:T},z)=f_{d}^{\mathsf{T}}(d^{1:T})f_{nxz}(n,x^{1:T},z)$. Here, similarly to classical ANOVA models, $f_{d}(\cdot)$ is a one-hot encoding vector for the entire (exponentially sized) space of possible $d_{n}^{1:T}$ trajectories, and $f_{nxz}(\cdot,\cdot,\cdot)$ returns a vector of corresponding length.
@@
As we do not want to restrict ourselves to fixed $T$ (and $N$), not to mention an exponential cost of representation, we will resort to fixed-dimensional embeddings of sequences and identities. In particular, we design $f_{d}(d^{1:t})$ to have a finite output dimensionality $r_{g}$. Inspired by sequential hidden representation models such as recurrent neural networks, we define $f_{dl}(d^{1:t}):=g_{l}(d^{1:t},t,t)$, for $l\in 1,2,\dots,r_{g}$, via the following recursive definition on $0\leq t^{\prime}\leq t$:
$g_{l}(d^{1:t^{\prime}},t^{\prime},t):=h_{l}(g_{l}(d^{1:t^{\prime}-1},t^{\prime}-1,t),d^{t^{\prime}},t^{\prime},t),$ (4)
@@
such that $h_{l}(v,0,t^{\prime},t)\equiv v$ for any $(v,t^{\prime},t)$ (that is, intervention level $D_{n}^{t^{\prime}}=0$ does nothing) and $h_{l}(v,d,0,t)\equiv 1$ for any $(d,v,t)$ (base case). A representational choice $h_{l}(v,d,t^{\prime},t):=h_{l_{1}}(v)^{\mathsf{T}}h_{l_{2}}(d,t^{\prime},t)$ requires a choice of dimensionality $r_{h_{l}}$ for the inner vectors $h_{l_{1}},h_{l_{2}}$. Recursively applying (4) in this product form can still be done relatively efficiently, but choosing $r_{h_{l}}=1$ for all $l$ simplifies matters as we apply again the inner product formula to $f_{d}^{\mathsf{T}}(d^{1:T})f_{nxz}(n,x^{1:T},z)$ to get to a factorization form analogous to Eq. (1). More formally:
@@
Let $f(n,d^{1:T},x^{1:T},z):=f_{d}^{\mathsf{T}}(d^{1:T})f_{nxz}(n,x^{1:T},z)$, where function sequences $f_{d}(d^{1:T})$ and $f_{nxz}(n,x^{1:T},z)$ are defined for all $T\in\mathbb{N}^{+}$ and have codomain $\mathbb{R}^{r_{g}}$. Furthermore, define $f_{dl}(d^{1:T})$ as in Eq. (4) with $r_{h}=1$, and the $l$-th entry of $f_{nxz}(n,x^{1:T},z)$ as $f_{nxzl}(n,x^{1:T},z):=u_{l}(n)^{\mathsf{T}}v_{l}(x^{1:T},z)$, each of the latter having codomain $\mathbb{R}^{r_{s^{\prime}}}$ and where $v_{l}(\cdot,\cdot)$ is defined recursively for any $T$, similarly $g_{l}(\cdot,\cdot,\cdot)$. Then there exists some integer $r$ and three functions $a,b$ and $c$ with codomain $\mathbb{R}^{r}$ so that $f(n,d^{1:T},x^{1:T},z)=\sum_{l=1}^{r}a_{l}(n)\times b_{l}(x^{1:T},z)\times c_{l}(d^{1:T},T)$. Moreover $c_{l}(d^{1:T},T)=\prod_{t=1}^{T}m_{l}(d^{t},t,T)$ for functions $m_{l}:\mathbb{N}^{3}\rightarrow\mathbb{R}$.
Proofs of this and next results are given in Appendix B. The above shows that if we start with an assumption that the ground truth function takes some desired form, then there is some finite $r_{g}$ for which our combination of basis functions can parameterize the ground truth function for arbitrary $T$, analogous to in Eq. (1). On the other hand, we show a companion result below, that if we hold $T$ fixed, there is some fixed value for $r_{g}$, which depends polynomially on $T$, such that any measurable ground truth function can be approximated by our combination of some appropriate choice of basis functions. This is reminiscent of the line of work on matrix factorization (e.g., Appendix A of Agarwal et al. (2023c)).
@@
Given a fixed value of $T$, suppose we have a measurable function $f:\bigcup_{t=2}^{T}\mathcal{X}^{1:t-1}\times\bigcup_{t=1}^{T}\mathcal{D}^{t}\times\mathcal{Z}\rightarrow\mathbb{R}^{d}$, $t\leq T$, where $\mathcal{X}^{1:t-1}$ and $\mathcal{Z}$ are compact, $\mathcal{D}^{t}$ is finite. For $d=1,\cdots,d_{max}$, let $m_{d}:\bigcup_{t=1}^{T}\mathcal{D}^{t}\rightarrow\{1,\cdots,T\}$ with $m_{d}(\mathbf{d})=t\text{ if }\exists t\;s.t.\;\mathbf{d}_{t}=d\;\text{else }0$. For every $T$ there exist $r_{g}=O(T^{d_{max}+1})$ such that for any $\epsilon$, there are measurable functions $\phi_{l}:\mathcal{X}^{1:t-1}\times\mathcal{Z}\rightarrow\mathbb{R}^{d}$ and $\psi_{ld^{\prime}}:\mathcal{D}^{t}\rightarrow\mathbb{R}$ and real numbers $\beta_{l}$ for $l=1,\cdots,r_{g}$ and $d=1,\cdots,d_{max}$, $\hat{f}(\mathbf{x},\mathbf{d},z)=\sum_{l=1}^{r_{g}}\phi_{l}(\mathbf{x},z)\beta_{l}\prod_{d=1}^{d_{max}}\psi_{ld^{\prime}}(m_{d^{\prime}}(\mathbf{d}))$ such that
$\text{sup}_{\mathbf{x}\in\bigcup_{t=2}^{T}\mathcal{X}^{1:t-1},\mathbf{d}\in\bigcup_{t=1}^{T}\mathcal{D}^{t},z\in\mathcal{Z}}\left|f(\mathbf{x},\mathbf{d},z)-\hat{f}(\mathbf{x},\mathbf{d},z)\right|\leq\epsilon.$ (5)
@@
Eq. (1) and Eq. (2) are motivated by the success of approaches for representation learning of sequences, which assumes that we can compile all the necessary information from the past using a current finite representation at any time step $t$. However, knowledge of Propositions 1 and 2 further imply that, for $f(n,d^{1:T},x^{1:T},z):=\mathbb{E}[X_{n}^{t}(d_{n}^{1:t})~{}|~{}x_{n}(d_{n}^{1:t-1}),z_{n}]$, the representation of functions in a multilinear combination of adaptive factors (Eq. (1)) is motivated by more fundamental results in functional analysis. Our assumption that no value of $d$ (other than the baseline 0) can be used more than once has mostly an empirical motivation, as this would imply unrealistic assumptions about the same intervention having the same effect when applied multiple times. Ultimately, we assume that the factors $h_{l_{2}}(d,t^{\prime},t)$ leading to Eq. (2) are time-translation invariant, being a function of $d$ and $t-t^{\prime}$ only. While it is possible to generalize beyond translation invariant models, as commonly done in the synthetic controls literature (Abadie, 2021) that inspired the causal matrix factorization ideas (Athey et al., 2021; Agarwal et al., 2023b, c), this would complicate matters further by requiring a fourth factor into the summation term of Eq. (1) to encode absolute time. Models for time-series forecasting with parameter drifts can be tapped into and integrated with our model family, but we leave these out as future work.
Interpretation.
At the starting of the process, where all units are assumed to be at their “natural” state (assuming $D_{n}^{1}=0$), we can interpret $\eta_{n}^{1}:=\beta_{n}$ as a finite basis representation, with respect to $\phi_{n}^{1}$, of $\mathbb{E}[X_{n}^{1}(0)~{}|~{}Z_{n}]=(\phi_{n}^{1})^{\mathsf{T}}\eta_{n}^{1}$. Each $\eta_{nl}^{1}$ can be seen as a latent feature of unit $n$. Effect vector $\psi_{n}^{t}$ defines a warping of $\eta_{n}^{t-1}$ that describes how treatments affect the behaviour of an individual with the resulting $\eta_{n}^{t}:=\eta_{n}^{t-1}\odot\psi_{n}^{t}$. The modifier $\psi_{n}^{t}$ can be interpreted as reverting, suppressing or promoting particular latent features that are assumed to encode all information necessary to reconstruct the conditional expectation of $X_{n}^{t}(d^{1:t})$ from the chosen basis (notice that intervention level 0 implies $\eta_{n}^{t}=\eta_{n}^{t-1}$ since we define $\psi_{l}(0,t^{\prime},t)=1$). This presents a sequential warping view of compositionality of interventions.
@@ 2.2 Identification and Data Assumptions
Assume for now that functions $\phi$ are known. As common in the literature of matrix factorization, assume also that we have access to some moments of the distribution. In particular, for given $N$ data points and $T$ time points, we have access to $\mathbb{E}[X_{n}^{t}(d_{n}^{1:t})~{}|~{}x_{n}^{t-1}(d_{n}^{1:t-1}),z_{n}]$. We will impose conditions on the realizations of $X_{n}^{1:T}$ and values chosen for $d_{n}^{1:T}$, as well as values of $N$ and $T$ as a function of $r$, in order to identify each $\beta_{n}$ and the parameters of $\psi_{l}(d,t^{\prime},t)$ for all intervention levels $d$ of interest. The theoretical results presented, which describe how parameters are identifiable from population expectations and known function spaces determined by a prescribed basis $\phi$ of known and finite dimensionality $r$, will then provide a basis for a practical learning algorithm in the sequel.
@@
let $A_{n}$ be a $(T_{0_{n}}-1)\times r$ matrix where each row $t=1,2,\dots,T_{0_{n}}-1$ is given by $\phi_{n}(x_{n}^{1:t}(0^{1:t}),z_{n})^{\mathsf{T}}$. We assume that $A_{n}$ is full rank with left pseudo-inverse $A_{n}^{+}$.
@@
Let $b_{n}$ be a $(T_{0_{n}}-1)\times 1$ vector with entries $\mathbb{E}[X_{n}^{t+1}(d_{n}^{1:{t+1}})~{}|~{}x_{n}^{1:t}(d_{n}^{1:t}),z_{n}]$. Under Assumptions 1 and 2, $\beta_{n}$ is identifiable from $A_{n}$ and $b_{n}$.
@@
Let $n_{1},n_{2},\dots,n_{N_{d}}$ index the elements of a set of units $\mathcal{N}_{d}$. For $t^{\prime}=1,2,\dots,k_{d}-1$, let $A_{dt^{\prime}}$ be a $N_{d}\times r$ matrix where each row $i=1,\dots,N_{d}$ is given by $\phi_{n_{i}}(x_{n_{i}}^{1:t+t^{\prime}-1}(d_{n_{i}}^{1:t+t^{\prime}-1}),z_{n_{i}})^{\mathsf{T}}$. We assume that each $A_{dt^{\prime}}$ is full rank with left pseudo-inverse $A_{dt^{\prime}}^{+}$.
@@
Assume we have a dataset of $N$ units and $T$ time points $(d_{n}^{1:T},x_{n}^{1:T},z_{n})$ generated by a model partially specified by Eq. (1). Assume also knowledge of the conditional expectations $\mathbb{E}[X_{n}^{t}(d_{n}^{1:t})~{}|~{}x_{n}^{t-1}(d_{n}^{1:t-1}),z_{n}]$ and basis functions $\phi_{l}(\cdot,\cdot)$ for all $l=1,2,\dots,r$ and all $1\leq n\leq N$ and $1\leq t\leq T$. Then, under Assumptions 1-4 applied to all individuals and all intervention levels $d$ that appear in our dataset, we have that all $\beta_{n}$ and all $\psi_{l}(d,\cdot,\cdot)$ are identifiable.
Notice that nothing above requires prior knowledge of all intervention levels which will exist, and could be applied on a rolling basis as new interventions are invented. There is no need for all units to start synchronously at $t=1$: the framing of the theory assumes so to simplify presentation, but the only requirement is that each unit is given a “burn-in” period of at least $T_{0}$ steps and that each new intervention level $d$ is applied to at least $r$ units which have not been perturbed recently by relatively novel interventions (that is, units $\mathcal{N}_{d}$ which, if they were given a past intervention not given to anyone else before, then that perturbation should have taken place enough steps in the past so that all recent but past $\psi$ functions have been identified).

### v2

@@ 2.1 Representation Power: Compositional Recursion and and Compositional Warping
Eq. (1) is reminiscent of tensor factorization and its uses in causal modeling (Athey et al., 2021; Agarwal et al., 2023b), where the primary motivation comes from the imputation of missing potential outcomes (Imbens and Rubin, 2015) taking place on a pre-determined period in the past. From a predictive perspective, it can be motivated from known results in functional analysis (such as Proposition 1 of (Kaddour et al., 2021)) that allows us to represent a function $f(x)$ by first partitioning its input into $(x_{a},x_{b})$ and controlling the approximation error given by an inner product of vector-valued functions,
$f(x_{a},x_{b})\approx f_{a}^{\mathsf{T}}(x_{a})f_{b}(x_{b}),$ (4)
which we call the factorization trick. In practice, we choose the dimensionality of such vectors $\{f_{a}(\cdot)$, $f_{b}(\cdot)\}$ by a data-driven approach. Although those results apply typically to smooth functions with continuous inputs (such as the Taylor series approximation, which relies on vectors of monomials with coefficients constructed from derivatives), discrete inputs such as $D_{n}^{1:T}$, at a fixed dimensionality $T$, can be separated from the other inputs as
$f(n,d^{1:T},x^{1:T},z):=f_{d}^{\mathsf{T}}(d^{1:T})f_{nxz}(n,x^{1:T},z).$ (5)
Here, similarly to classical ANOVA models, $f_{d}(\cdot)$ is a one-hot encoding vector for the entire (exponentially sized) space of possible $d_{n}^{1:T}$ trajectories. That is, $f_{d}(\cdot)$ is binary with exactly one non-zero entry, corresponding to which of the possible trajectories $d_{n}^{1:T}$ took place. The codomain of $f_{nxz}(\cdot,\cdot,\cdot)$ must be a set of vectors of corresponding (exponential in $T$) length.
In what follows, we will start from the premise of further partitioning $f(n,d^{1:T},x^{1:T},z):=\mathbb{E}[X_{n}^{T+1}~{}|~{}x_{n}^{1:T},z_{n},\mathrm{do}(d_{n}^{1:T})]$ by applying the factorization trick of Eq. (4) repeatedly, justifying the generality of the right-hand side (RHS) of Eq. (1). We will accomplish this by first introducing more structure into the respective $f_{d}(d^{1:T})$, based on a recursive composition of functions. We will then show in Proposition 1 how the RHS of Eq. (1) follows. We also provide a complementary theoretical result in Proposition 2 which, by assuming a bounded $T$, describes the choice of dimensionality in the application of Eq. (4) as controlling a formal approximation error for a general class of functions.
Main results: factorization trick with compositional recursion.
@@
As we do not want to restrict ourselves to fixed $T$ (and $N$), not to mention an exponential cost of representation, we will resort to fixed-dimensionality embeddings of sequences and identities. In particular, we design $f_{d}(d^{1:t})$ to have a finite and tractable output dimensionality $r_{g}$ for all $t$. Inspired by sequential hidden representation models such as recurrent neural networks, for each entry $l\in 1,2,\dots,r_{g}$ in the output of $f_{d}(\cdot)$, we define $f_{dl}(d^{1:t})$ via the following recursive definition:
$\begin{array}[]{rcl}f_{dl}(d^{1:t})&:=&g_{l}(d^{1:t},t,t),\\ g_{l}(d^{1:t^{\prime}},t^{\prime},t)&:=&h_{l}(g_{l}(d^{1:t^{\prime}-1},t^{\prime}-1,t),d^{t^{\prime}},t^{\prime},t),\hskip 14.45377pt0<t^{\prime}\leq t,\\ h_{l}(v,d,t^{\prime},t)&:=&h_{l_{1}}(v)^{\mathsf{T}}h_{l_{2}}(d,t^{\prime},t).\\ \end{array}$ (6)
The base case for $t^{\prime}=0$ is $g_{l}(d^{1:0},0,t)\equiv 1$ for all $t$, with $d^{1:0}$ undefined. Moreover, we define $h_{l}(v,0,t^{\prime},t)\equiv v$ for any $(v,t^{\prime},t)$ – that is, intervention level $D_{n}^{t^{\prime}}=0$ does nothing.
@@
In order to define $h_{l}(v,d,t^{\prime},t)$ for $d>0$, we will apply the same trick of splitting the input into two subsets, and defining a vector representation on each. The representational choice $h_{l}(v,d,t^{\prime},t):=h_{l_{1}}(v)^{\mathsf{T}}h_{l_{2}}(d,t^{\prime},t)$ requires a choice of dimensionality $r_{h_{l}}$ for the inner vectors $h_{l_{1}},h_{l_{2}}$. Recursively applying (6) in this product form can still be done relatively efficiently, but choosing $r_{h_{l}}=1$ for all $l$ simplifies matters. Finally, we apply the factorization trick one last time to define $f_{nxz}(n,x^{1:T},z)$ within Eq. (5) to get to a factorization form analogous to Eq. (1). More formally:
@@
Let (i) $f(n,d^{1:T},x^{1:T},z):=f_{d}^{\mathsf{T}}(d^{1:T})f_{nxz}(n,x^{1:T},z)$, where function sequences $f_{d}(d^{1:T})$ and $f_{nxz}(n,x^{1:T},z)$ are defined for all $T\in\mathbb{N}^{+}$ and have codomain $\mathbb{R}^{r_{g}}$; (ii) $f_{dl}(d^{1:T})$ be given as in Eq. (6) with $r_{h_{l}}=1$; (iii) the $l$-th entry of $f_{nxz}(n,x^{1:T},z)$ be given by $f_{nxzl}(n,x^{1:T},z):=u_{l}(n)^{\mathsf{T}}v_{l}(x^{1:T},z)$, where both $u_{l}(\cdot)$ and $v_{l}(\cdot,\cdot)$ have codomain $\mathbb{R}^{r_{s^{\prime}}}$; (iv) $v_{l}(x^{1:t},z)$ be defined analogously to $f_{dl}(d^{1:t})$ as in Eq. (6), carrying out the fixed-size $z$ as an extra argument. Then there exists some integer $r$ and three functions $a,b$ and $c$ with codomain $\mathbb{R}^{r}$ so that $f(n,d^{1:T},x^{1:T},z)=\sum_{l=1}^{r}a_{l}(x^{1:T},z)\times b_{l}(n)\times c_{l}(d^{1:T},T)$. Moreover $c_{l}(d^{1:T},T)=\prod_{t=1}^{T}m_{l}(d^{t},t,T)$ for some function $m_{l}:\mathbb{N}^{3}\rightarrow\mathbb{R}$.
Proofs of this and next results are given in Appendix B. The above shows that if we start with an assumption that the ground truth function takes some desired form, then there is some finite $r$ for which our combination of basis functions can parameterize the ground truth function for arbitrary $T$, analogous to Eq. (1), where $(a,b,c)$ plays the role of $(\phi,\beta,\psi)$. On the other hand, we show a companion result below: that if we hold $T$ fixed, there is some fixed value for $r$, which depends polynomially on $T$, such that any measurable ground truth function can be approximated by our combination with some appropriate choice of basis functions.
@@
Given a fixed value of $T$, suppose we have a measurable function $f:\bigcup_{t=2}^{T}\mathcal{X}^{1:t-1}\times\bigcup_{t=1}^{T}\mathcal{D}^{t}\times\mathcal{Z}\rightarrow\mathbb{R}^{d}$, $t\leq T$, where $\mathcal{X}^{1:t-1}$ and $\mathcal{Z}$ are compact, $\mathcal{D}^{t}$ is finite. Let $\mathbf{x}\in\bigcup_{t=2}^{T}\mathcal{X}^{1:t-1}$ and let $\mathbf{d}\in\bigcup_{t=1}^{T}\mathcal{D}^{t}$. For $d=1,\cdots,d_{max}$, let $m_{d}:\bigcup_{t=1}^{T}\mathcal{D}^{t}\rightarrow\{1,\cdots,T\}$ with $m_{d}(\mathbf{d})=t\text{ if }\exists t\;s.t.\;\mathbf{d}_{t}=d\;\text{else }0$. Then, for any $\epsilon>0$, there exist $r=O(T^{d_{max}+1})$ and measurable functions $\phi_{l}:\mathcal{X}^{1:t-1}\times\mathcal{Z}\rightarrow\mathbb{R}^{d}$ and $\psi_{ld^{\prime}}:\mathcal{D}^{t}\rightarrow\mathbb{R}$, and real numbers $\beta_{l}$, such that
$\text{sup}_{\mathbf{x}\in\bigcup_{t=2}^{T}\mathcal{X}^{1:t-1},\mathbf{d}\in\bigcup_{t=1}^{T}\mathcal{D}^{t},z\in\mathcal{Z}}\left|f(\mathbf{x},\mathbf{d},z)-\sum_{l=1}^{r}\phi_{l}(\mathbf{x},z)\times\beta_{l}\times\prod_{d=1}^{d_{max}}\psi_{ld^{\prime}}(m_{d^{\prime}}(\mathbf{d}))\right|\leq\epsilon.$ (7)
@@
The theoretical results above are reminiscent of the line of work on matrix factorization (e.g., Appendix A of Agarwal et al. (2023c)). Eq. (1) and Eq. (2) are motivated by the success of approaches for representation learning of sequences, which assume that we can compile all the necessary information from the past using a current finite representation at any time step $t$. However, Propositions 1 and 2 further imply that, for $f(n,d^{1:T},x^{1:T},z):=\mathbb{E}[X_{n}^{T}~{}|~{}x_{n}^{1:T-1},z_{n},\mathrm{do}(d_{n}^{1:T})]$, the representation of functions in a multilinear combination of adaptive factors (Eq. (1)) is motivated by more fundamental results in functional analysis. Our assumption that no value of $d$ (other than the baseline 0) can be used more than once has mostly an empirical motivation, as this would imply unrealistic assumptions about the same intervention having the same effect when applied multiple times. Ultimately, we assume that the factors $h_{l_{2}}(d,t^{\prime},t)$ leading to Eq. (2) are time-translation invariant, being a function of $d$ and $t-t^{\prime}$ only. While it is possible to generalize beyond translation invariant models, as done in the synthetic controls literature (Abadie, 2021) that inspired causal matrix factorization ideas (Athey et al., 2021; Agarwal et al., 2023b, c), this would complicate matters further by requiring a fourth factor into the summation term of Eq. (1) encoding absolute time. Models for time-series forecasting with parameter drifts can be tapped into and integrated with our model family, but we leave these out as future work.
Interpretation: compositional warping.
At the start of the process, where all units are assumed to be at their “natural” state (assuming $D_{n}^{1}=0$), we can interpret $\eta_{n}^{1}:=\beta_{n}$ as a finite basis representation, with respect to $\phi_{n}^{1}$, of $\mathbb{E}[X_{n}^{1}~{}|~{}Z_{n},\mathrm{do}(d_{n}^{1}=0)]=(\phi_{n}^{1})^{\mathsf{T}}\eta_{n}^{1}$. Each $\eta_{nl}^{1}$ can be seen as a latent feature of unit $n$. Effect vector $\psi_{n}^{t}$ defines a warping of $\eta_{n}^{t-1}$ that describes how treatments affect the behaviour of an individual in feature space, with the resulting $\eta_{n}^{t}:=\eta_{n}^{t-1}\odot\psi_{n}^{t}$. The modifier $\psi_{n}^{t}$ can be interpreted as reverting, suppressing or promoting particular latent features that are assumed to encode all information necessary to reconstruct the conditional expectation of $X_{n}^{t}$ from the chosen basis (notice that intervention level 0 implies $\eta_{n}^{t}=\eta_{n}^{t-1}$ since we define $\psi_{l}(0,t^{\prime},t)=1$). This presents a sequential warping view of compositionality of interventions. Our framing also conveniently reduces the problem of identifiability of conditional effects to the problem of identifying baseline vectors $\beta_{n}$ and warp function embeddings $\psi_{l}(d,t^{\prime},t)$, as we will see next.
@@ 2.2 Identification and Data Assumptions
Assume for now that functions $\phi$ are known. As commonly done in the matrix factorization literature, assume also that we have access to some moments of the distribution. In particular, for given $N$ data points and $T$ time points, we have access to $\mathbb{E}[X_{n}^{t}~{}|~{}x_{n}^{1:t-1},z_{n},\mathrm{do}(d_{n}^{1:t})]$. We will impose conditions on the realizations of $X_{n}^{1:T}$ and values chosen for $d_{n}^{1:T}$, as well as values of $N$ and $T$ as a function of $r$, in order to identify each $\beta_{n}$ and the parameters of $\psi_{l}(d,t^{\prime},t)$ for all intervention levels $d$ of interest. The theoretical results presented, which describe how parameters are identifiable from population expectations and known function spaces determined by a prescribed basis $\phi$ of known and finite dimensionality $r$, will then provide the foundation for a practical learning algorithm in the sequel.
@@
Let $A_{n}$ be a $(T_{0_{n}}-1)\times r$ matrix, each row $t=1,2,\dots,T_{0_{n}}-1$ is given by $\phi_{n}(x_{n}^{1:t},z_{n})^{\mathsf{T}}$ realizations under regime $\mathrm{do}(d_{n}^{1:t})$. Let $A_{n}$ be full rank with left pseudo-inverse $A_{n}^{+}$.
The first assumption can be interpreted as allowing for a “burn-in” period for unit $n$ under only the default action. The second assumption ensures enough diversity among realized features $\phi_{n}$ so that a least-squares projection can be invoked to identify $\beta_{n}$, as formalized in the following proposition.
@@
Let $b_{n}$ be a $(T_{0_{n}}-1)\times 1$ vector with entries $\mathbb{E}[X_{n}^{t+1}~{}|~{}x_{n}^{1:t},z_{n},\mathrm{do}(d_{n}^{1:{t+1}})]$. Under Assumptions 1 and 2, $\beta_{n}$ is identifiable from $A_{n}$ and $b_{n}$.
The next assumption is the requirement that any action level $d$ of interest is carried out across a large enough number of units. Moreover, they must be sufficiently separated in time from follow-up non-default action levels so that their parameters can be identified. Assumption 4 is a counterpart to Assumption 2, ensuring linear independence of feature trajectories.
@@
Let $n_{1},n_{2},\dots,n_{N_{d}}$ index the elements of a set of units $\mathcal{N}_{d}$. For $t^{\prime}=1,2,\dots,k_{d}-1$, let $A_{dt^{\prime}}$ be a $N_{d}\times r$ matrix where each row $i=1,\dots,N_{d}$ is given by $\phi_{n_{i}}(x_{n_{i}}^{1:t+t^{\prime}-1},z_{n_{i}})^{\mathsf{T}}$ realized under regime $\mathrm{do}(d_{n_{i}}^{1:t+t^{\prime}-1})$. We assume that each $A_{dt^{\prime}}$ is full rank with left pseudo-inverse $A_{dt^{\prime}}^{+}$.
@@
Assume we have a dataset of $N$ units and $T$ time points $(d_{n}^{1:T},x_{n}^{1:T},z_{n})$ generated by a model partially specified by Eq. (1). Assume also knowledge of the conditional expectations $\mathbb{E}[X_{n}^{t}~{}|~{}x_{n}^{t-1},z_{n},\mathrm{do}(d_{n}^{1:t})]$ and basis functions $\phi_{l}(\cdot,\cdot)$ for all $l=1,2,\dots,r$ and all $1\leq n\leq N$ and $1\leq t\leq T$. Then, under Assumptions 1-4 applied to all individuals and all intervention levels $d$ that appear in our dataset, we have that all $\beta_{n}$ and all $\psi_{l}(d,\cdot,\cdot)$ are identifiable.
Notice that nothing above requires prior knowledge of all intervention levels which will exist, and could be applied on a rolling basis as new interventions are invented. There is no need for all units to start synchronously at $t=1$: the framing of the theory assumes so to simplify presentation, with the only requirement being that each unit is given a “burn-in” period of at least $T_{0}$ steps and that each new intervention level $d$ is applied to at least $r$ units which have not been perturbed recently by relatively novel interventions. Ultimately, the gap between a novel intervention level and the next one should be sufficiently large according to the complexity of the model, as indexed by $r$. This makes explicit that we do not get free identification: the more complex the domain requirements, the more information we need, both in terms of the time waited and the number of observations.